OpenStudy (anonymous):

Can someone please help me solve this pre cal question? \[\log_{12}(k-2) +7=9\] I got k=4 but I don't think I did it right..

OpenStudy (anonymous):

no solution

OpenStudy (anonymous):

can you tell me what steps you did to figure that out?

OpenStudy (anonymous):

log_12 (k-2)=9-7 log_12 (k-2)=2

OpenStudy (anonymous):

i also got k=146 s:

OpenStudy (anonymous):

so, (k-2)=12^2 k-2=144 k=146

OpenStudy (anonymous):

how did you get the 12^2?

OpenStudy (anonymous):

146, yeah ^_^

OpenStudy (anonymous):

oh i forgot about that rule!! thanks so much y'all :)

OpenStudy (anonymous):

you are welcome.....@ali1029